Income Statement- December 31, 2014 to December 31, 2015 Operating Revenues and Expenses Operating Revenues $1,000 Sales Returns 22,400 250 22,150 Total Operating Revenues Operating Expenses Cost of goods and services sold 6,000 3,000 6,000 900 250 6,000 600 22,750 Labor Materials Indirect cost Selling and promotion Depreciation General and administrative Lease payment Total Operating Expense Total Operating Income -600 Nonoperating Revenues and Expenses Rents Interest receipts Interest payments 30 200 -80 150 Total nonoperating Income Net Income before Tax Income Taxes Net Profit (loss) for 2015 -450 -450 Determine the following 1. Net Profit/Loss before tax 2. Net Profit/Loss Ratio 3. Interest Coverage 4. Identify the reasons behind the bad performance of this company.
Answer #1

1.Net Loss before tax=$-450

2.Net loss ratio= Net loss/ Net sales*100

=$ -450/ $22150*100=-2.03%

3.Interest Coverage ratio= EBIT/ Interest

EBIT= Net profit afer tax+ tax paid+ Interest paid- Interest received=

= -$450+$0+$370-$200=$ -280

=Interest coverage ratio=-280/200=-1.4

4.Company is running into losses. It is due to increased burden of operating and non operating expenses. Sales must be improves and effective measures to control the operating and non operating cost must be undertaken.